# A Novel State-Centric Necessary Condition for Time-Optimal Control of Controllable Linear Systems Based on Augmented Switching Laws (Extended Version)

### v2

@@ V-B Feasibility of the Perturbed BBS Trajectory
[…]
The feasibility of the perturbed trajectory is proved by two steps. Firstly, in Proposition 7, the state error between the perturbed trajectory and the original trajectory is bounded by $C\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|$ where $C$ is a constant dependent on the original trajectory. In other words, if the perturbation of keypoints’ time is sufficiently small and the augmented switching law is fixed, then the states of the perturbed trajectory are close to those of the original trajectory. Secondly, Theorem 2 proves the perturbed trajectory’s feasibility based on Propositions 4, 5, 6, and 7. In other words, if the original trajectory is feasible, the perturbation of keypoints’ time is sufficiently small, and the augmented switching law is fixed, then the perturbed trajectory is also feasible. Specifically, Proposition 7 helps to provide an upper bound on the perturbation under the requirement of the perturbed trajectory’s feasibility in Theorem 2.
@@
Proof.
Therefore, $\forall t\in\left[\hat{t}_{2i},\hat{t}_{2i+1}\right]$, it holds that
$\left\|{\boldsymbol{x}}\left(t\right)-{\boldsymbol{x}}^{\prime}\left(t\right)\right\|\leq\left\|{\mathrm{e}}^{{\boldsymbol{A}}_{i}\left(t-\hat{t}_{2i}\right)}\right\|\left\|{\boldsymbol{x}}\left(\hat{t}_{2i}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i}\right)\right\|$ (32)
$\leq$ $C_{3}\left\|{\boldsymbol{x}}\left(\hat{t}_{2i}\right)-{\boldsymbol{x}}^{\prime}\left(\hat{t}_{2i}\right)\right\|.$
$\forall i\in\left[M\right]$, let
$\begin{dcases}j=i+1,\,j^{\prime}=i,&\text{if }t_{i}\leq t_{i}^{\prime},\\ j=i,\,j^{\prime}=i+1,&\text{if }t_{i}>t_{i}^{\prime}.\end{dcases}$ (33)
Then, $\forall t\in\left[\hat{t}_{2i-1},\hat{t}_{2i}\right]$, $\dot{\boldsymbol{x}}={\boldsymbol{A}}_{j}{\boldsymbol{x}}+{\boldsymbol{b}}_{j}$ and $\dot{\boldsymbol{x}}^{\prime}={\boldsymbol{A}}_{j^{\prime}}{\boldsymbol{x}}+{\boldsymbol{b}}_{j^{\prime}}$. Lemma 1 implies that $\forall t\in\left[\hat{t}_{2i-1},\hat{t}_{2i}\right]$,
Therefore, (30) holds.
@@
Consider a feasible non-chattering BBS trajectory ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ whose ASL is ${\mathcal{L}}$ in problem (1a). Then, $\exists\varepsilon>0$ dependent on ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$, s.t. $\forall{\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ is trajectory also represented by ${\mathcal{L}}^{\prime}={\mathcal{L}}$, if $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$, then ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ is feasible for problem (1a). ${\boldsymbol{x}}\left(t\right)$ and ${\boldsymbol{x}}^{\prime}\left(t\right)$ are induced by ${\boldsymbol{t}}$ and ${\boldsymbol{t}}^{\prime}$ in the same way as that in Proposition 7.
Proof.
Fix ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$. Firstly, let
Secondly, consider the inequality constraints ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}+d_{p}\leq 0$. Denote $\left\{i^{(p)}_{j}\right\}_{j=0}^{N_{p}}\subset\left[M\right]\cup\left\{0\right\}$ increasing strictly monotonically, s.t. $i^{(p)}_{0}=0$, $i^{(p)}_{N_{p}}=M$, and $\forall j\in\left[N_{p}-1\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\leq 0$ switches between active and inactive at $t_{i^{(p)}_{j}}$. In other words, $\forall j\in\left[N_{p}-1\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t_{i_{j}^{(p)}}\right)+d_{p}=0$ holds. Furthermore, $\forall j\in\left[N_{p}\right]$, $t\in\left(t_{i^{(p)}_{j-1}},t_{i^{(p)}_{j}}\right)$, either ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}\equiv 0$ or ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}<0$ holds. Since ${\mathcal{L}}={\mathcal{L}}^{\prime}$, the above property on $\left({\boldsymbol{x}}\left(t\right),{\boldsymbol{t}}\right)$ holds for $\left({\boldsymbol{x}}^{\prime}\left(t\right),{\boldsymbol{t}}^{\prime}\right)$ as well.
Consider the case where ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}+d_{p}\equiv 0$ in $\left(t_{i^{(p)}_{j-1}},t_{i^{(p)}_{j}}\right)$, as shown in Fig. 3(a). ${\mathcal{L}}={\mathcal{L}}^{\prime}$ implies that ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}+d_{p}\equiv 0$ holds. Therefore, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}\leq 0$ is feasible in $\left(t_{i^{(p)}_{j-1}}^{\prime},t_{i^{(p)}_{j}}^{\prime}\right)$.
Denote $f_{p}\left(t\right)={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}$ and $f_{p}^{\prime}\left(t\right)={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}\left(t\right)+d_{p}$. Then, $\forall r\in{\mathbb{N}}^{*}$, $\frac{\mathrm{d}^{r}f_{p}}{\mathrm{d}t^{r}}={\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}_{i_{j}^{(p)}}^{r-1}\left({\boldsymbol{A}}_{i_{j}^{(p)}}{\boldsymbol{x}}+{\boldsymbol{b}}_{i_{j}^{(p)}}\right)$ is continuous in $\left(t_{i^{(p)}_{j-1}},t_{i^{(p)}_{j-1}+1}\right)$. If $f_{p}\left(t_{i^{(p)}_{j-1}}\right)<0$ holds, then let $r_{j}^{(p)}=0$ and $t_{1,j}^{(p)}=t_{i_{j-1}^{(p)}}$; otherwise, according to Propositions 4, 5, and 6, $\exists r_{j}^{(p)}\in\left[n\right]$, s.t. $\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)<0$ and $\forall r\in\left[r_{j}^{(p)}-1\right]$, $\frac{\mathrm{d}^{r}f_{p}}{\mathrm{d}t^{r}}\left(t_{i^{(p)}_{j-1}}^{+}\right)=0$. Let $t_{1,j}^{(p)}\in\left(t_{i^{(p)}_{j-1}},t_{i^{(p)}_{j-1}+1}\right)$ is sufficiently small, s.t. $\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t\right)<\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)<0$ holds $\forall t\in\left(t_{i^{(p)}_{j-1}},t_{1,j}^{(p)}\right)$. $t_{2,j}^{(p)}\leq t_{i^{(p)}_{j}}$ is constructed similarly. Then, $f_{p}\left(t\right)$ decreases strictly monotonically in $\left(t_{i^{(p)}_{j-1}},t_{1,j}^{(p)}\right)$ and increases strictly monotonically in $\left(t_{2,j}^{(p)},t_{i^{(p)}_{j}}\right)$.
For the case where $f_{p}\left(t_{i^{(p)}_{j-1}}\right)<0$ and $t_{1,j}^{(p)}=t_{i_{j-1}^{(p)}}$, let $\varepsilon$ be sufficiently small, s.t. $f_{p}\left(t\right)\leq\frac{1}{2}f_{p}\left(t_{i^{(p)}_{j-1}}\right)$ in $\left[t_{i^{(p)}_{j-1}}^{\prime},t_{i^{(p)}_{j-1}}\right]$. By (30) and (38), $\forall t\in\left[t_{i^{(p)}_{j-1}}^{\prime},t_{i^{(p)}_{j-1}}\right]$, $f_{p}^{\prime}\left(t\right)\leq 0$ is feasible since
$f_{p}^{\prime}\left(t\right)\leq$ $f_{p}\left(t\right)+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}$ (40)
$\leq$ $\frac{1}{2}f_{p}\left(t_{i^{(p)}_{j-1}}\right)-\frac{1}{4}f_{p}\left(t_{i^{(p)}_{j-1}}\right)<0.$
For the case where $f_{p}\left(t_{i^{(p)}_{j-1}}\right)=0$ and $t_{i_{j-1}^{(p)}}<t_{1,j}^{(p)}$, let
$\varepsilon<\hat{\varepsilon}\triangleq-\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)\left(4C\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{A}}_{i_{j}^{(p)}}\right\|^{r_{j}^{(p)}}+1\right)^{-1}.$ (41)
$\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t\right)\leq\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t\right)+\left\|{\boldsymbol{c}}_{p}\right\|\left\|{\boldsymbol{A}}_{i_{j}^{(p)}}\right\|^{r_{j}^{(p)}}\left\|{\boldsymbol{x}}-{\boldsymbol{x}}^{\prime}\right\|_{\infty}$ (42)
$\leq$ $\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)-\frac{1}{4}\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)<0.$
$\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t\right)\leq{\boldsymbol{c}}_{p}^{\top}{\boldsymbol{A}}_{i_{j}^{(p)}}^{r_{j}^{(p)}-1}\left({\boldsymbol{A}}_{i_{j}^{(p)}}{\boldsymbol{x}}\left(t_{i^{(p)}_{j-1}}\right)+{\boldsymbol{b}}_{i_{j}^{(p)}}\right)$ (44)
$\leq$ $\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)-\frac{1}{2}\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}}{\mathrm{d}t^{r_{j}^{(p)}}}\left(t_{i^{(p)}_{j-1}}^{+}\right)<0.$
Therefore, once $\varepsilon$ is sufficiently small, then $\frac{\mathrm{d}^{r_{j}^{(p)}}f_{p}^{\prime}}{\mathrm{d}t^{r_{j}^{(p)}}}<0$ holds in $\left[t_{i^{(p)}_{j-1}}^{\prime},t_{1,j}^{(p)}\right]$. ${\mathcal{L}}^{\prime}={\mathcal{L}}$ implies that $\forall r\in\left[r_{j}^{(p)}-1\right]$, $\frac{\mathrm{d}^{r}f_{p}^{\prime}}{\mathrm{d}t^{r}}\left(t_{i^{(p)}_{j-1}}^{+}\right)=0$; hence, $f_{p}^{\prime}$ decreases strictly monotonically in $\left(t_{i^{(p)}_{j-1}}^{\prime},t_{1,j}^{(p)}\right)$. Therefore, $\forall t\in\left[t_{i^{(p)}_{j-1}}^{\prime},t_{1,j}^{(p)}\right]$, ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}<0$ is feasible. Similarly, once $\varepsilon$ is sufficiently small, then ${\boldsymbol{c}}_{p}^{\top}{\boldsymbol{x}}^{\prime}\left(t\right)+d_{p}<0$ is feasible in $\left[t_{2,j}^{(p)},t_{i^{(p)}_{j}}^{\prime}\right]$.
@@ V-C State-Centric Necessary Condition of Optimality
Proof.
$\forall i>1$, $j<i-1$, $\frac{\partial{\boldsymbol{x}}_{i}}{\partial t_{j}}=\prod_{k=i}^{j+2}{\mathrm{e}}^{{\boldsymbol{A}}_{k}\left(t_{k}-t_{k-1}\right)}\frac{\partial{\boldsymbol{x}}_{j+1}}{\partial t_{j}}$ holds, which implies (45). ∎
@@ VI Applications in Chain-of-Integrator Systems
$\min\quad$ $J=\int_{0}^{{t_{\mathrm{f}}}}\mathrm{d}t,$ (48a)
[…]
@@
[…]
@@ VI-A The Case where $\forall k\in\left[n\right]$, $x_{{\mathrm{m}}k}=\infty$
Proof.
Chattering does not occur since no state inequality constraints exist. Assume that the optimal control switches $\left(N-1\right)$ times. In other words, $\forall i\in\left[N\right]$, $t\in\left(t_{i-1},t_{i}\right)$, $u\left(t\right)\equiv u_{i}\in\left\{\pm{u_{\mathrm{m}}}\right\}$. $u_{i}\not=u_{i-1}$, implies that $\Delta u_{i}\not=0$. Then,
${\boldsymbol{J}}=\left[-\Delta u_{i+1}{\boldsymbol{\phi}}\left(t_{N}-t_{i}\right)\right]_{i=1}^{N-1}\in{\mathbb{R}}^{n\times\left(N-1\right)}.$ (49)
The Vandermonde form of ${\boldsymbol{J}}$ implies that ${\mathrm{rank}}{\boldsymbol{J}}=\min\left\{n,N-1\right\}$. Theorem 3 implies that ${\mathrm{rank}}{\boldsymbol{J}}<n$; hence, $N-1<n$. ∎
@@ VI-B The Case where Only System Behaviors Occur
Proof.
For a constrained arc ${\mathcal{S}}_{i}$, add the equality constraint at $t_{i}$, i.e., ${\boldsymbol{x}}_{i,1:\left|{\mathcal{S}}_{i}\right|}={\mathrm{sgn}}\left({\mathcal{S}}_{i}\right)x_{{\mathrm{m}}\left|{\mathcal{S}}_{i}\right|}{\boldsymbol{e}}_{\left|{\mathcal{S}}_{i}\right|}$. Note that if ${\mathcal{S}}_{i}$ is constrained, i.e., $\left|{\mathcal{S}}_{i}\right|\not=0$, then $u\equiv 0$ in ${\mathcal{S}}_{i}$. Therefore, the constraints induced by ${\mathcal{L}}$ are
$\begin{dcases}{\boldsymbol{x}}_{i,1:\left|{\mathcal{S}}_{i}\right|}={\mathrm{sgn}}\left({\mathcal{S}}_{i}\right)x_{{\mathrm{m}}\left|{\mathcal{S}}_{i}\right|}{\boldsymbol{e}}_{\left|{\mathcal{S}}_{i}\right|},\,\forall i\in\left[N\right],\,\left|{\mathcal{S}}_{i}\right|\not=0,\\ {\boldsymbol{x}}_{N}={{\boldsymbol{x}}_{\mathrm{f}}}.\end{dcases}$ (52)
By Condition 3, $\left|{\mathcal{S}}_{N}\right|=0$. Denote $\left|{\mathcal{S}}_{N+1}\right|=n$, and =$\mathcal{I}=\left\{i\in\left[N+1\right]:\left|{\mathcal{S}}_{i}\right|\not=0\right\}$. According to Proposition 8, ${\boldsymbol{J}}$ contains the following rows:
${\boldsymbol{J}}=\left(-\Delta u_{j+1}{\boldsymbol{\phi}}_{1:\left|{\mathcal{S}}_{i}\right|}\left(t_{i}-t_{j}\right)\delta_{j<i}\right)_{i\in\mathcal{I},\,j\in\left[N-1\right]},$ (53)
where $\delta_{A}$ is the indicator function of condition $A$, i.e.,
$\delta_{A}=\begin{dcases}1,&\text{condition }A\text{ holds},\\ 0,&\text{condition }A\text{ does not hold}.\end{dcases}$ (54)
Let ${\boldsymbol{A}}\leftrightarrow{\boldsymbol{B}}$ mean that ${\boldsymbol{A}}$ has full row rank if and only if ${\boldsymbol{B}}$ has full row rank. By $\forall j\in\left[N-1\right]$, $\Delta u_{j+1}\not=0$, it holds that
${\boldsymbol{J}}\leftrightarrow\left({\boldsymbol{\phi}}_{1:\left|{\mathcal{S}}_{i}\right|}\left(t_{i}-t_{j}\right)\delta_{j<i}\right)_{i\in\mathcal{I},\,j\in\left[N-1\right]}.$ (55)
By iteratively taking differences between adjacent rows of (55), dividing by the time difference, and eliminating, it can be proved that ${\boldsymbol{J}}$ has full rank based on Conditions 1, 2, and 3. The proof is similar to our previous work [20] and is omitted due to space limitation. According to Theorem 3, the number of rows should be greater than the number of columns, i.e.,
$\sum_{i\in\mathcal{I}}\left|{\mathcal{S}}_{i}\right|=\sum_{i=1}^{N}\left|{\mathcal{S}}_{i}\right|+n>N-1.$ (56)
Therefore, (50) holds. ∎
@@
Proof.
Since $\forall k\in{\mathbb{N}}^{*}$, ${\boldsymbol{x}}_{1:2}\left(t_{3k-3}\right)={\boldsymbol{x}}_{1:2}\left(t_{3k}\right)=x_{{\mathrm{m}}2}{\boldsymbol{e}}_{2}$, it can be proved that the control $u$ in $\left(t_{3k-3},t_{3k}\right)$ is
$\begin{dcases}u\left(t\right)\equiv u_{3k-2}=-{u_{\mathrm{m}}},&t\in\left(t_{3k-3},t_{3k-2}\right),\\ u\left(t\right)\equiv u_{3k-1}={u_{\mathrm{m}}},&t\in\left(t_{3k-2},t_{3k-1}\right),\\ u\left(t\right)\equiv u_{3k}=-{u_{\mathrm{m}}},&t\in\left(t_{3k-1},t_{3k}\right).\end{dcases}$ (60)
Among them, $t_{3k-1}=\frac{3t_{3k}+t_{3k-3}}{4}$ and $t_{3k-2}=\frac{t_{3k}+3t_{3k-3}}{4}$.
$\forall N\in{\mathbb{N}}^{*}$, the ASL between ${\boldsymbol{x}}\left(t_{0}\right)$ and ${\boldsymbol{x}}\left(t_{3N}\right)$ is denoted as ${\mathcal{L}}_{N}$, which requires that $\forall k\in\left[N-1\right]$, ${\boldsymbol{x}}_{3k,1:2}=x_{{\mathrm{m}}2}{\boldsymbol{e}}_{2}$, and ${\boldsymbol{x}}_{3N}={\boldsymbol{x}}\left(t_{3N}\right)$. By Proposition 8, the Jacobian matrix ${\boldsymbol{J}}$, except the last column, is (51). Through basic row and column transformations, it holds that ${\boldsymbol{J}}\leftrightarrow{\boldsymbol{J}}^{\prime}$. Considering the rank of columns $\left(k-n+3\right)$ through $k$ of ${\boldsymbol{J}}^{\prime}$, (59) holds.
@@
However, it is challenging to obtain the recursive equation (59) for problem (48a) based on the traditional PMP-based necessary condition [9] since the behaviors of costates are significantly complex for high-order problems, let alone determining the existence of chattering. A comparison between state-centric and PMP-based approaches is provided in Appendix A.

### v3

@@ V-B Feasibility of the Perturbed BBS Trajectory
Consider a feasible BBS trajectory ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ of problem (1a) without chattering. The ASL is denoted as $\mathcal{L}$}. For convenience of discussion, the keypoints of $\mathcal{L}$} occur at $t_{i}$, where $0=t_{0}<t_{1}<\dots<t_{M}=t_{\mathrm{f}}$. In other words, $\left\{t_{i}\right\}_{i=0}^{M}$ consists of time points at ends of arcs and tangent markers. Denote ${\boldsymbol{t}}=\left(t_{i}\right)_{i=1}^{M}$, and $\dot{\boldsymbol{x}}={\boldsymbol{A}}_{i}{\boldsymbol{x}}+{\boldsymbol{b}}_{i}$ in $\left(t_{i-1},t_{i}\right)$. This section discusses the feasibility of a perturbed trajectory ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ represented by the ASL $\mathcal{L}$$^{\prime}$, where the time vector ${\boldsymbol{t}}^{\prime}=\left(t_{i}^{\prime}\right)_{i=1}^{M}$ is close to $\boldsymbol{t}$}, i.e., $\left\|{\boldsymbol{t}}^{\prime}-{\boldsymbol{t}}\right\|$ is sufficiently small. For convenience, the notation $\left\|\bullet\right\|$ refers to $\left\|\bullet\right\|_{\infty}$ in this section if not specified.
The feasibility of the perturbed trajectory is proved by two steps. Firstly, in Proposition 7, the state error between the perturbed trajectory and the original trajectory is bounded by $C\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|$ where $C$ is a constant dependent on the original trajectory. In other words, if the perturbation of keypoints’ time is sufficiently small and the augmented switching law is fixed, then the states of the perturbed trajectory are close to those of the original trajectory. Secondly, Theorem 2 guarantees the feasibility of the perturbed trajectory based on Propositions 4, 5, 6, and 7. In other words, if the original trajectory is feasible, the perturbation of keypoints’ time is sufficiently small, and the augmented switching law is fixed, then the perturbed trajectory is also feasible. Specifically, Proposition 7 helps to provide an upper bound on the perturbation under the requirement of the perturbed trajectory’s feasibility in Theorem 2.
@@
Consider a feasible non-chattering BBS trajectory ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$ whose ASL is $\mathcal{L}$} in problem (1a). Then, $\exists\varepsilon>0$ dependent on ${\boldsymbol{x}}={\boldsymbol{x}}\left(t\right)$, s.t. $\forall{\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ is a trajectory also represented by $\mathcal{L}$$^{\prime}$$\mathcal{L}$}, if $\left\|{\boldsymbol{t}}-{\boldsymbol{t}}^{\prime}\right\|<\varepsilon$, then ${\boldsymbol{x}}={\boldsymbol{x}}^{\prime}\left(t\right)$ is feasible for problem (1a). ${\boldsymbol{x}}\left(t\right)$ and ${\boldsymbol{x}}^{\prime}\left(t\right)$ are induced by $\boldsymbol{t}$} and $\boldsymbol{t}$$^{\prime}$ in the same way as that in Proposition 7.
@@ VI Applications in Chain-of-Integrator Systems
$\min\quad$ ${t_{\mathrm{f}}}=\int_{0}^{{t_{\mathrm{f}}}}\mathrm{d}t,$ (26a)
${\mathrm{s.t.}}\quad$ $\dot{x}_{k}\left(t\right)=x_{k-1}\left(t\right),\,1<k\leq n,\,\forall t\in\left[0,t_{\mathrm{f}}\right],$ (26b)
@@
$\left|x_{k}\left(t\right)\right|\leq x_{{\mathrm{m}}k},\,1\leq k\leq n,\,\forall t\in\left[0,t_{\mathrm{f}}\right],$ (26e)
@@ VI-B The Case where Only System Behaviors Occur
However, it is challenging to obtain the recursive equation (30) for problem (26a) based on the traditional PMP-based necessary condition [9] since the behaviors of costates are significantly complex for high-order problems, let alone determining the existence of chattering. A comparison between state-centric and PMP-based approaches is provided in Appendix B.